Penicillin is hydrolyzed and thereby rendered inactive by penicillinase (also known as β-lactamase), an enzyme present in some resistant bacteria. The amount of penicillin hydrolyzed in 1 minute in a 10-ml solution containing 10-13 moles of purified penicillinase was measured as a function of the concentration of penicillin. Assume that the concentration of penicillin does not change appreciably during the assay.

I got the right answers for the first two parts but I can't figure out how to find the turnover number. Please give a step by step for the last part! Thank you
Answer #1

From the graph we can say that it follows Michaleis-Menton Kinetics.

9000 y = 7.6234x + 1.4638 in N/T 7 N O 0.2 0.4 0.8 1 1.2 0.6 1/[S]

Now, y=1/V

x= 1/S

for x=0, y=1.4638. Let this value of y be called as A.

A=1/Vmax = 1.4638

Thus, Vmax = 1/A = 1/1.4638 = 0.68315

For the equation y=mx+c, m is the slope.

Hence, Slope=7.6234

Now, Slope= KM / Vmax

So, 7.6234= KM / 0.68315

KM = 7.6234*0.68315 = 5.2079

Now, Turnover No. = Vmax / Total enzyme concentration

Here, Enzyme concentration = 10-13 moles

Turnover no. = 0.68315/ 10-13

= 683.15*1010